On May 10, 2015, Substantial Corporation spent $210,000 on a limited-life intangible asset. Its useful lifetime is 10 years. How much in total amortization expense should be recorded on the intangible asset by December 31, 2017?

$ -0-
$42,000
$56,000
$63,000
Answer #1

Answer

  • Correct Answer = Option #3 $ 56,000
  • Annual amortisation for 12 months = $ 210,000 / 10 years = $ 21,000
  • 2015’s amortisation = $ 21000 x 8 months/12 months = $ 14,000 [May to Dec = 8 months]
    2016’s amortisation = $ 21000
    2017’s amortisation = $ 21,000
  • Total amortisation by Dec 31 2017 = 14000 + 21000 + 21000 = $ 56,000
  • Answer = $ 56000